OpenStudy (anonymous):

Any help would be very much appreciated:3) A third staircase in a swimming pool has a landing of 15 inches and a vertical rise of 12 inches per step. To the nearest degree, calculate the following measures (you must show all work to receive full credit): a) What is the angle of elevation rounded to the nearest whole degree? b) If the elevation of each step were changed to 10 inches and the landing became 7 inches, what is the new angle of elevation for the staircase to the nearest whole degree?

OpenStudy (anonymous):

Yes I copied and pasted it but, here is what I know... the angle elevation is 55 not sure how to show it though.

OpenStudy (kohai):

 Let angle of elevation be θ tan θ = 12/15 θ = tan⁻¹ (12/15) = ?

OpenStudy (kohai):

Ugh let me draw this

OpenStudy (kohai):

 Do you know why I use inverse tan instead of regular tan?